If the second term is 1/2 and the seventh term of a geometric progression is 1/64 find the 10th term
Answer by stanbon(75887)

If the second term is 1/2 and the seventh term of a geometric progression is 1/64 find the 10th term
-----
2nd term:: ar = 1/2
7th term:: ar^6 = 1/64
-----
Divide to solve for "r":
r^5 = 2^-5
r = 2^-1
r = 1/2
-----
Solve for "a"
ar = 1/2
a(1/2) = 1/2
a = 1
-------
Ans: a(n) = (1/2)^(n-1)
So, a(10) = (1/2)^9 = 1/512
=================================
Cheers,
Stan H.
============